# Jakub Poćwiardowski, Mateusz Modrzejewski

### v2

@@ Abstract
This work introduces the $\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ composer system, capable of generating complete, multi-minute musical compositions with complex structures in any time signature, in the MIDI domain from input descriptions in natural language. The system utilizes an autoregressive transformer language model to map natural language prompts to composition parameters in JSON format. The defined structure includes time signature, scales, chord progressions, and valence-arousal values, from which accompaniment, melody, bass, motif, and percussion tracks are created. We propose a genetic algorithm for the generation of melodic elements. The algorithm incorporates mutations with musical significance and a fitness function based on normal distribution and predefined musical feature values. The values adaptively evolve, influenced by emotional parameters and distinct playing styles. The system for generating percussion in any time signature utilises probabilistic methods, including Markov chains. Through both human and objective evaluations, we demonstrate that our music generation approach outperforms baselines on specific, musically meaningful metrics, offering a viable alternative to purely neural network-based systems.
@@ 1 Introduction
All the current state-of-the-art methods use neural networks to effectively replicate human music composition abilities. However, they depend on broad datasets, leading to reliance on dominant structures like 4/4 meter and common chord progressions. Furthermore, text-to-music models often fail to respond to specific musical terms, and audio-based music generation models produce outputs that are hard to edit further.
In this work, we present $\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$, a tool for generating full multi-track songs in symbolic format and their subsequent editing with further prompts without relying on dominant musical structures. The system uses an autoregressive transformer model to map the input text to generated song structure and attributes. Further, it utilizes genetic algorithms along with probabilistic methods for generating MIDI (Musical Instrument Digital Interface) tracks for specific sections.
@@ 2.1 Large language models for music generation
The recent revolution of Large Language Models (LLMs) invoked experiments concerning their abilities in tasks they were not initially created for. Examples of LLM usage for music composition include ChatMusician [10] which is a fine-tuned Llama2 model to treat Music ABC notation as a language, and ComposerX [11] which uses a multi-agent approach and chain-of-thought, where distinct instances of GPT-4 take care of interpreting user inputs, creating melodies, harmonising, and choosing instruments, among others.
@@ 3 Structure of $M^{6}(GPT)^{3}$
The structure of the system is presented in fig. 1. Although there is no direct feedback to the LLM, the generated structure and previous prompts provide context that allows for further modification.
@@ 3.1 Predicting structure from text
To generate song structure and parameters, a LLM is used. We chose the GPT family because of its promising music theory knowledge and ability to use it via openai API, making it accessible from any computer, contrary to locally hosted LLMs requiring high computing resources.
@@
The complete JSON template [^1] Full template and system prompt available at https://jpocwiar.github.io/M6-GPT3-Composer-Demo/. contains:
@@ 4 Evolutionary Melodic Track Generator
    Melodies - the most unrestricted tracks, playing mostly in mid-range, but capable of reaching high and low notes.
@@ 4.3 Fitness Function
When designing the fitness function, we adapted the approach from [15], which employs a Gaussian distribution. In our implementation, the distribution’s mean represents the target metric value. The melody’s fitness score is determined by the weighted sum of each metric’s distribution value at the point corresponding to the actual measured metric value of the melody. The metrics used for genetic algorithm optimization include: the mean percentage of unique notes and intervals per measure; percentages of dissonant intervals, intervals larger than an octave, notes in the scale, notes in the current chord and rests in a sequence; tonal range; mean percentage of unique note lengths per measure; average note pitch and its deviation; length of strong beat notes; melodic contour; off-beat notes; normalized average interval size; logarithm of average note length and its deviation; counts of consecutive intervals within (1,3) and short notes (eighth or shorter); length of repeated fragments; and the number of measures beginning with a root note. All measured values are normalized to the (0, 1) range. The formula for fitness function calculation is presented in eq. 1.
@@
To ensure harmonic coherence in a multitrack composition, we add another metric to assess dissonance between distinct tracks. Our scoring method, based on [18], addresses these issues and also accounts for rests in one of the tracks. We evaluate intervals between distinct tracks modulo 12 at each index using scores from table 2. The total score is computed as the average of these individual scores. Finally, we normalize the average score for the entire song $a_{e}$ using the $tanh$ function as detailed in eq. 2.
@@ 4.4 Melodic Tracks Modes
To support the diverse characteristics of melodic sections and assist LLM decision-making, we define three melodic sections and 11 generation modes, which influence genetic algorithm parameters.
@@
Bass tracks include four modes: Short riff (one-measure riff repeated and shifted to fit chord roots), Long riff (two-measure riff with similar behavior), Bassline (lowest chord roots with randomized transitions and mutations for emotion), and Repetitive bassline (like bassline, but with consistent mutations per bar).
Motifs offer five modes: Long motif (one measure, repeated and shifted), Opening motif (half-measure with rests in the second half), Closing motif (rests in the first half), Repeated motif (half-measure repeated), and Short repeated motif (quarter-note length, repeated and trimmed to fit the sequence).
@@ 5 Percussion Generator
To generate songs in odd time signatures, Deep Learning models like Transformers may not be ideal. Our analysis of the well-known Groove dataset [21] revealed that 98.96% of drum loops are in 4/4 time. Given the relative simplicity and commonality of drum patterns in many songs, we opted for a rule-based system incorporating probability and Markov Chains.
@@ 5.1 Data Representation
5.2 Bass Drum and Snare
For the bass drum and snare, we analyzed fundamental patterns from a variety of sources, including drum courses, publicly available resources, and drum-specific datasets. Based on this analysis, we manually constructed a table of probabilities for these components in time signatures built on quarter-note and eighth-note divisions, with the number of beats per bar ranging from 2 to 9 (e.g., time signatures from 2/4 to 9/8).
For time signatures not covered by the predefined patterns, we apply a recursive decomposition approach. For instance, a 13/8 time signature is broken down into 7/8 and 6/8 patterns, while a 25/4 signature is split into 13/4 and 12/4, and further into 7/4, 6/4, 6/4, and 6/4. For time signatures with shorter notes, such as 11/16, we use patterns for longer notes (11/8 in that case).
As previously mentioned, each drum state is encoded with a sixteenth-note interval. Due to this representation, we pad the patterns with silent states, as detailed in eq. 3. Specifically, time signatures based on quarter notes include three silent states for each beat, while those based on eighth notes include one silent state per beat, and so forth.
$I=\frac{16}{{M_{R}}}-1$ (3)
where:
  • $I$ is a number of zero states added after every beat,
  • $M_{R}$ is the type of note (lower numeral of the time signature).
Hi-hats accent the tempo and are placed on quarter, eighth, or sixteenth notes, with patterns varying by the section’s emotion. Open hi-hats occasionally replace the last closed hi-hat in a measure to mark transitions, influenced by the arousal parameter. Ride cymbals and bells typically play on quarter notes, offset by an eighth, with variations based on emotional cues. Crash cymbals primarily occur at the start of measures, rarely appearing elsewhere.
Drum fills use all five General MIDI toms and the snare, generated based on fill probability and length. Fills are constructed using Markov chains, where each state is a 5-bit sequence representing drum activations, ordered from lowest to highest tom.
After adding all previously mentioned drum components, we obtain a full list of 12-bit sequences. To enhance diversity and realism, several modifications are applied to the patterns. The bass drum and snare drum occurrences are repeated in the following states with a certain probability. For states where the sum of active bits representing elements played with hands (bits 1, 2, and 5–12) exceeds 2, bits in this range are randomly zeroed until the sum is reduced to 2. Additionally, all bits from 5 to 12 are randomly zeroed with a specified probability for each state.
@@ 6 Chord generator
Chord progressions are the foundation of the composition, defined for each section based on the language model’s responses. The chord track is generated in one of three modes. In Continuous mode, the chord notes are sustained throughout, creating a seamless background. In Repeated mode, the notes are played together with lengths and rests determined by the arousal parameter, with up to 40% repetitions randomly omitted. In Arpeggio mode, the notes are played sequentially up and down the scale, with durations ranging from a sixteenth to a half note depending on arousal.
The chord names provided by LLM do not specify the octaves in which the individual notes are to be played, nor the possible repetition of these notes in other octaves. For $\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ Composer, we use valence-arousal values to apply chord inversions and modify voicing sizes. We partially base the solution on [16], where the authors propose that at the minimum value of arousal, a chord should consist of two notes, and at the maximum - of eight notes. In the context of our multi-track system, it was decided to limit the maximum number of notes in a chord to six. To ensure that the system works well for chords with varying numbers of notes (e.g., seventh or ninth chords), the following adjustments are made: If the arousal is below 0.3 or the chord has more than 4 notes, a fifth of the chord is removed. If the arousal exceeds 0.7 and the chord has fewer than 5 notes, a note an octave above the root is added. If the arousal exceeds 0.9 and the chord has fewer than 6 notes, a note forming an octave and the chord’s fifth interval is added.
Concerning the octaves in which the chord notes are played, [16] proposes concentrating the notes around the C4 note for the minimum value of valence and C6 for the maximum value. Taking into account that for our system, chords serve only as an accompaniment for a multitrack piece, with melody playing in a similar pitch register, it was decided to shift the range of central note concentration to span from A2 to A4. Based on the value of valence, the center of gravity of the notes is determined. Then, for each chord, the appropriate overturns are made using optimization to ensure that the notes do not overlap and that their center of gravity is as close as possible to the desired location. The conditioning of the chords via valence and arousal is presented in table 4.
@@ 7.2 Subjective Listening Test
The listening test was divided into two segments. In the first segment (table 6), participants evaluated 15 compositions from 3 different systems. Each composition was rated on a 1-to-5 scale across five criteria: 1) richness and diversity, 2) memorability, 3) entertainment value, 4) emotional conveyance, and 5) inspirational quality. The comparison systems were ComposerX [11] and MMT [9]. We selected 5 "best examples" that were more than a minute in length from the websites of both systems, randomly mixed them with our model’s 5 best compositions, and presented them to participants for evaluation.
Surveyed qualities (1-5 scale)
RD M EV CE I Avg
@@
Table 5: Comparison against the baseline models (RD: Richness and diversity, M: Memorability, EV: Entertainment value, CE: Emotional conveyance, I: Inspirational quality, Avg: Average of all criteria). Mean values and 95% confidence intervals are reported. According to the listeners, $\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ outperformed the other evaluated models in 4 out of 5 criteria, with its most significant advantage being the entertainment value of its compositions.
In the second test (LABEL:tab:subjective-results-midicaps), we compared the music generated by our system with pieces from the recently released MIDICaps dataset [23], which currently is the only openly available large-scale MIDI dataset with text captions. We randomly selected 5 full-length pieces from the dataset and generated 5 corresponding full-length songs using our system based on the same descriptions [^4] Generated pieces can be heard on https://jpocwiar.github.io/M6-GPT3-Composer-Demo/.. The participants were then surveyed to determine whether both the dataset compositions and the $\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ generations accurately matched their descriptions regarding tempo, instruments, mood, structure, and overall impression.
Surveyed qualities (1-5 scale)
RD M EV CE I Avg
Table 6: Comparison against the baseline models (RD: Richness and diversity, M: Memorability, EV: Entertainment value, CE: Emotional conveyance, I: Inspirational quality, Avg: Average of all criteria). Mean values and 95% confidence intervals are reported. According to the listeners, $\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ outperformed the other evaluated models in 4 out of 5 criteria, with its most significant advantage being the entertainment value of its compositions.
@@ 7.3 Objective Evaluation
To further support the listening test, we follow [9] and evaluate pitch class entropy, scale consistency, and groove consistency using the MusPy framework [24, 25]. For our system, we assess these metrics in three different scenarios. The “Prompt-driven” scenario involves generating full songs from text using an LLM. The “Focused” scenario standardizes the structure with a 4/4 time signature, 120 bpm tempo, C-F-Am-F chord progression in C Major and middle emotional values across all generations. In the “Randomized” scenario, all parameters are random except for the chords, which belong to the same scale. We compare our results with those presented in [9] and display them in table 7. We find that the outputs of $\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ have comparable musical traits to a reference set of real music.
Objective evaluation results
@@
$\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ Prompt-driven (ours) $\mathbf{2.907}\pm 0.067$ $\mathbf{92.640\pm 2.11}$ $98.919\pm 0.15$
$\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ Focused (ours) $2.849\pm 0.010$ $98.804\pm 0.14$ $98.558\pm 0.01$
$\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ Randomized (ours) $2.886\pm 0.021$ $95.113\pm 0.94$ $98.890\pm 0.11$
Table 7: Comparison with the baseline models with objective metrics following [9]. Mean values and 95% confidence intervals are reported, with the closest values to the ground truth highlighted in bold. Across three established metrics, $\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ with LLM-conditioned song structure achieves the best performance in two out of three cases.
@@ 8 Conclusions
In this work, we introduce $\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$, a method for generating music from text without relying on extensive music datasets. Both objective metrics and subjective studies show that this approach can produce realistic and inspiring compositions. We believe that as fully neural-based systems continue to evolve, there is an increasing need to emphasize optimization-based techniques to create innovative and creative tools. The combination of controllability and structural flexibility of our approach with neural advances such as style morphing has the potential to push the boundaries of music generation in the future.

### v3

@@ Abstract
This work introduces the M6(GPT)3 composer system, capable of generating complete, multi-minute musical compositions with complex structures in any time signature, in the MIDI domain from input descriptions in natural language. The system utilizes an autoregressive transformer language model to map natural language prompts to composition parameters in JSON format. The defined structure includes time signature, scales, chord progressions, and valence-arousal values, from which accompaniment, melody, bass, motif, and percussion tracks are created. We propose a genetic algorithm for the generation of melodic elements. The algorithm incorporates mutations with musical significance and a fitness function based on normal distribution and predefined musical feature values. The values adaptively evolve, influenced by emotional parameters and distinct playing styles. The system for generating percussion in any time signature utilises probabilistic methods, including Markov chains. Through both human and objective evaluations, we demonstrate that our music generation approach outperforms baselines on specific, musically meaningful metrics, offering a viable alternative to purely neural network-based systems.
@@ 1 Introduction
All the current state-of-the-art methods use neural networks to effectively replicate human music composition abilities. However, they depend on broad datasets, leading to reliance on dominant structures like 4/4 meter and common chord progressions. Furthermore, text-to-music models often fail to respond to specific musical terms, and audio-based music generation models produce outputs that are hard to edit further. To address these limitations, we introduce the M6(GPT)3 system, designed for flexible multi-track song generation and editing in symbolic format.
Our contributions may be summarized as follows:
  • We propose M6(GPT)3, a novel solution for generating and editing full multi-track songs in symbolic format without relying on dominant musical structures. We integrate genetic algorithms and probabilistic methods to generate MIDI tracks for specific song sections.
  • We evaluate M6(GPT)3 using subjective and objective, musically meaningful metrics.
  • We release demo examples and Python code to facilitate reproducibility and further research.
@@ 2.1 Large language models for music generation
The recent revolution of Large Language Models (LLMs) invoked experiments concerning their abilities in tasks they were not initially created for. Examples of LLM usage for music composition include ComposerX [10] which uses a multi-agent approach and chain-of-thought, where distinct instances of GPT-4 take care of interpreting user inputs, creating melodies, harmonising, and choosing instruments.
@@ 3 Structure of M6(GPT)3
The system structure is shown in fig.˜1. While the LLM receives no direct feedback, the generated output and prior prompts provide context for further modification.
@@ 3.1 Predicting structure from text
To generate song structure and parameters, an LLM is used. We chose the GPT family because of its promising music theory knowledge and ability to use it via openai API, making it accessible from any computer, contrary to locally hosted LLMs requiring high computing resources. Note that the system’s architecture features an interchangeable LLM component, enabling the integration of more advanced models.
@@
The complete JSON template contains:
@@ 4 Evolutionary Melodic Track Generator
    Melodies - the most unrestricted tracks, mostly mid-range but capable of high and low notes.
@@ 4.3 Fitness Function
When designing the fitness function, we adapted the approach from [14], which employs a Gaussian distribution. In our implementation, the distribution’s mean represents the target metric value. The melody’s fitness score is determined by the weighted sum of each metric’s distribution value at the point corresponding to the actual measured metric value of the melody. The metrics used for genetic algorithm optimization include: the mean percentage of unique notes and intervals per measure; percentages of dissonant intervals, intervals larger than an octave, notes in the scale, notes in the current chord and rests in a sequence; tonal range; mean percentage of unique note lengths per measure; average note pitch and its deviation; length of strong beat notes; melodic contour; off-beat notes; normalized average interval size; logarithm of average note length and its deviation; counts of consecutive intervals within (1,3) and short notes (eighth or shorter); length of repeated fragments; and the number of measures beginning with a root note. All measured values are normalized to the (0, 1) range. The formula for fitness function calculation is presented in eq.˜1.
@@
To ensure harmonic coherence in multitrack compositions, we introduce a metric to assess inter-track dissonance. Based on [17], our method accounts for rests and evaluates intervals modulo 12 using table˜2. The total score is the average of these values, normalized via the tanhtanh function as in eq.˜2.
@@ 4.4 Melodic Tracks Modes
To support the diverse characteristics of melodic sections and assist LLM decision-making, we arbitrarily define three melodic sections and 11 generation modes, which influence genetic algorithm parameters.
@@
Bass tracks include four modes: Short riff (one-measure riff repeated and shifted to fit chord roots), Long riff (two-measure riff with similar behavior), Bassline (lowest chord roots with randomized transitions and mutations for emotion), and Repetitive bassline (bassline with same mutations per bar).
Motifs offer five modes: Long motif (one measure, repeated and shifted), Opening motif (half-measure with rests in the second half), Closing motif (rests in the first half), Repeated motif (half-measure repeated), and Short repeated motif (quarter-note length, repeated and trimmed).
@@ 5 Percussion Generator
To generate songs in odd time signatures, Deep Learning models like Transformers may not be ideal, as training datasets are heavily biased toward 4/4. Our analysis of the well-known Groove dataset [20] revealed that 98.96% of its drum loops are in 4/4 time. Given the relative simplicity and repetitiveness of drum patterns in many songs, we opted for a rule-based system incorporating probability and Markov Chains.
@@ 5.1 Data Representation
5.2 Drum Patterns
We analyzed fundamental drum patterns from various sources to construct probability tables for bass drum and snare placement in time signatures from 2/4 to 9/8. For unsupported signatures, we use recursive decomposition (e.g., 13/8 splits into 7/8 and 6/8), while shorter-note signatures (e.g., 11/16) inherit patterns from longer equivalents (e.g., 11/8). Each drum state is encoded in sixteenth-note intervals, requiring silent padding: three silent states per beat for quarter-note-based signatures, one for eighth-note-based signatures, and so forth.
Hi-hats emphasize tempo, appearing on quarter, eighth, or sixteenth notes, with patterns shaped by the section’s emotion. Open hi-hats often replace the final closed hi-hat in a measure to mark transitions, influenced by the arousal parameter. Ride cymbals and bells typically follow quarter-note pulses, offset by an eighth, with emotional variations. Crash cymbals primarily occur at measure starts, rarely appearing elsewhere.
5.3 Toms and Drum Fills
Drum fills are generated using Markov chains, guided by fill probability and length. Each state represents a 5-bit pattern for tom activations, with a separate state for the snare.
After assembling the 12-bit drum sequences, we enhance diversity and realism by modifying patterns. Bass and snare hits may repeat probabilistically. If hand-played elements (bits 1, 2, 5–12) exceed two in a state, random bits are zeroed to reduce the count. Additionally, bits 5–12 are randomly zeroed per state with a set probability.
@@ 6 Chord generator
Chord progressions form the composition’s foundation, defined per section by the language model. The chord track is generated in one of three modes: Continuous (sustained chords), Repeated (notes played with arousal-dependent lengths, omitting up to 40% of repetitions), or Arpeggio (notes played sequentially with durations from a sixteenth to a half note based on arousal).
Since the LLM does not specify octaves or voicing, we adjust inversions and note counts using valence-arousal values, partially following [15]. Chord size varies from two notes (low arousal) to six (high arousal), with adjustments: removing the fifth for low arousal or large chords, adding an octave-above note for high arousal if the chord is small, and including both an octave and fifth at extreme arousal levels.
For octave placement, we modify [15]’s approach to fit our multitrack system, shifting the note concentration range from A2 to A4 to avoid overlap with the melody. The center of gravity is determined by valence, and chords are optimized to maintain spacing and minimize overlap. Conditioning via valence and arousal is detailed in table˜4.
@@ 7.2 Subjective Listening Test
The listening test was divided into two segments. In the first segment (table˜5), participants evaluated 15 compositions from 3 different systems. Each composition was rated on a 1-to-5 scale across five criteria: 1) richness and diversity, 2) memorability, 3) entertainment value, 4) emotional conveyance, and 5) inspirational quality. The comparison systems were ComposerX [10] and MMT [9]. We selected 5 "best examples" that were more than a minute in length from the websites of both systems, randomly mixed them with our model’s 5 best compositions, and presented them to participants for evaluation.
RD M EV EC I Avg
@@
Table 5: Comparison against the baseline models (RD: Richness and diversity, M: Memorability, EV: Entertainment value, CE: Emotional conveyance, I: Inspirational quality, Avg: Average of all criteria). Mean values and 95% confidence intervals are reported. According to the listeners, M6(GPT)3 outperformed the other evaluated models in 4 out of 5 criteria, with its most significant advantage being the entertainment value of its compositions.
Tempo Instruments Mood Structure Overall
M6(GPT)3 generation 4.06 ± 0.28 4.25 ± 0.24 3.58 ± 0.32 3.61 ± 0.41 3.80 ± 0.25
Actual described song from MIDICaps 4.14 ± 0.21 4.38 ± 0.17 3.79 ± 0.23 4.09 ± 0.26 4.06 ± 0.18
Table 6: Comparison of original song from MIDICaps dataset and M6(GPT)3 generation from the same description evaluating how well each matches that description. Mean values and 95% confidence intervals are reported. While the generations produced by M6(GPT)3 generally underperform compared to the actual described songs, their results demonstrate a notable proximity to the original compositions.
In the second test (table˜6), we compared our system’s outputs to MIDICaps [22], the only large-scale public MIDI dataset with text descriptions. We randomly selected 5 full-length pieces from the dataset and generated 5 corresponding full-length songs using our system based on the same descriptions.  [^3] Generated pieces can be heard on https://jpocwiar.github.io/M6-GPT3-Composer-Demo/. The participants were then surveyed to determine whether both the dataset compositions and the M6(GPT)3 generations accurately matched their descriptions regarding tempo, instruments, mood, structure, and overall impression.
@@ 7.3 Objective Evaluation
To further support the listening test, we follow [9] and evaluate pitch class entropy, scale consistency, and groove consistency using the MusPy framework [23, 24]. For our system, we assess these metrics in three different scenarios. The “Prompt-driven” scenario involves generating full songs from text using an LLM. The “Focused” scenario standardizes the structure with a 4/4 time signature, 120 bpm tempo, C-F-Am-F chord progression in C Major and middle emotional values across all generations. In the “Randomized” scenario, all parameters are random except for the chords, which belong to the same scale. We compare our results with those presented in [9] and display them in table˜7. We find that the outputs of $\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ have comparable musical traits to a reference set of real music.
@@
$\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ - Prompt-driven (ours) $\mathbf{2.907}\pm 0.067$ $\mathbf{92.640\pm 2.11}$ $98.919\pm 0.15$
$\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ - Focused (ours) $2.849\pm 0.010$ $98.804\pm 0.14$ $98.558\pm 0.01$
$\text{M}^{\text{6}}(\text{GPT})^{\text{3}}$ - Randomized (ours) $2.886\pm 0.021$ $95.113\pm 0.94$ $98.890\pm 0.11$
Table 7: Comparison with the baseline models with objective metrics following [9]. Mean values and 95% confidence intervals are reported, with the closest values to the ground truth highlighted in bold. Across three established metrics, M6(GPT)3 with LLM-conditioned song structure achieves the best performance in two out of three cases.
@@ 8 Conclusions
In this work, we introduce M6(GPT)3, a novel method for generating music from text without relying on extensive music datasets. Both objective metrics and subjective studies show that this approach can produce realistic and inspiring compositions. To facilitate further research, we publish code and generated examples. We believe that as fully neural-based systems continue to evolve, there is an increasing need to emphasize optimization-based techniques to create innovative and creative tools. The combination of controllability and structural flexibility of our approach with neural advances such as style morphing has the potential to push the boundaries of music generation in the future.